OpenStudy (anonymous):

The world's highest fountain of water is located, appropriately enough, in Fountain Hills, Arizona. The fountain rises to a height of 560 ft. (a) What is the initial speed of the water? (b) How long does it take for the water to reach the top of the fountain?

OpenStudy (shane_b):

height (h) = 560ft acceleration due to gravity (g) = 32 ft/s^2 To find the initial velocity start with the kinematic equation: \[V_f^2-V_i^2=2ad\]Rearrange it (g=a and h=d) to get:\[\Delta V=\sqrt{2gh}\]Plug in known values:\[\Delta V=\sqrt{2(32ft/s^2)(560ft)}=189.3ft/s\]To find the time it takes to get to the top:\[t=\frac{\Delta V}{g}=\frac{189.3ft/s}{32ft/s^2}=5.9s\]

OpenStudy (anonymous):

Thank you.

OpenStudy (shane_b):

yw